OpenStudy (clarence):

How would I go abouts solving a 3x4 matrix system?

OpenStudy (clarence):

\[\left[\begin{matrix}5 & 3 & -9 & 2 \\ -4 & -4 & -8 & 5 \\ 21 & 11 & -53 & 15\end{matrix}\right]\]

OpenStudy (clarence):

Specifically one such as above, except with a line down the middle before the last column?

ganeshie8 (ganeshie8):

\[\left[\begin{matrix}5 & 3 & -9 &|& 2 \\ -4 & -4 & -8 &|& 5 \\ 21 & 11 & -53 &|& 15\end{matrix}\right]\] like this ?

OpenStudy (clarence):

Yes, how did you manage to do that on here by the way?

ganeshie8 (ganeshie8):

you may right click and see latex commands : 

OpenStudy (clarence):

Oh, okay then.

ganeshie8 (ganeshie8):

btw, we call it a 3x3 system the size of matrix \(A\) is \(3\times 3\) in the system of equations : \[Ax=b\]

ganeshie8 (ganeshie8):

you're just attaching that "b" also as the last column to get that 3x4 matrix that 3x4 matrix with "b" attached is called "augmented matrix"

ganeshie8 (ganeshie8):

familiar with elimination/row opearations ?

OpenStudy (clarence):

Familiar but not well enough to solve the problem at hand unfortunately.

ganeshie8 (ganeshie8):

first see that the solutions are not affected by row operations because you're just adding/subtracting a multiple of one row from the other... this doesn't change the solutions

ganeshie8 (ganeshie8):

if you have x + y = 2 x+2y = 3 then, adding one equation to other won't change the solutions because you're adding the same thing to both sides of a particular equation

ganeshie8 (ganeshie8):

first convince yourself that below two systems will have same solutions : system1 ``` x + y = 2 x+2y = 3 ``` system2 ``` x + y = 2 2x+3y = 5 ```

OpenStudy (clarence):

Okay..

ganeshie8 (ganeshie8):

your goal in elimination is to get a triangular matrix

OpenStudy (clarence):

So, not 4x + 6y = 10?

ganeshie8 (ganeshie8):

that is a valid row operation too, you're allowed to multiply entire row by a constant

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

you want to change the starting matrix to above form

ganeshie8 (ganeshie8):

start by working the element at 2,1 \[\left[\begin{matrix}5 & 3 & -9 &|& 2 \\ -4 & -4 & -8 &|& 5 \\ 21 & 11 & -53 &|& 15\end{matrix}\right]\] adding 4/5 times first row to second row does the job, right ?

OpenStudy (clarence):

Yes, at least I think so anyway

ganeshie8 (ganeshie8):

do it

OpenStudy (clarence):

So now the second and third row match?

ganeshie8 (ganeshie8):

no, what do you get after doing that first step

OpenStudy (clarence):

\[\left[\begin{matrix}5 & 3 & -9 & | 2 \\ 21 & 11 & -53 & |15 \\ 21 & 11 & -53 & |15\end{matrix}\right]\]

ganeshie8 (ganeshie8):

how ?

OpenStudy (clarence):

5 times the 1st row added into the 2nd row?

ganeshie8 (ganeshie8):

yeah you could do that!

OpenStudy (clarence):

Was I not supposed to do that?

ganeshie8 (ganeshie8):

that is not a standard method, but what you did is a better one..

ganeshie8 (ganeshie8):

so lets keep going

ganeshie8 (ganeshie8):

for second step, maybe subtract second row from third row

ganeshie8 (ganeshie8):

\[\left[\begin{matrix}5 & 3 & -9 & | 2 \\ 21 & 11 & -53 & |15 \\ 21 & 11 & -53 & |15\end{matrix}\right]\] R3-R2 gives : \[\left[\begin{matrix}5 & 3 & -9 & | 2 \\ 21 & 11 & -53 & |15 \\ 0& 0& 0& |0\end{matrix}\right]\]

OpenStudy (clarence):

Okay then, what's the next step?

ganeshie8 (ganeshie8):

next look at the element at 2,1 position you want to make it 0 so that you get a triangular matrix

OpenStudy (clarence):

So 21 needs to be 0?

ganeshie8 (ganeshie8):

yes how to do that using row operations ?

OpenStudy (clarence):

Multiply the 2nd row by 5 and then add 21x1st row to the second row?

ganeshie8 (ganeshie8):

Multiply the 2nd row by 5 and then "subtract" 21x1st row to the second row?

ganeshie8 (ganeshie8):

you mean that right ?

OpenStudy (clarence):

Yes, sorry I was thinking in terms of -21

ganeshie8 (ganeshie8):

that works perfectly! do it

OpenStudy (clarence):

\[\left[\begin{matrix}5 & 3 & -9 & |2 \\ 0 & -8 & -76 & |33 \\ 0 & 0 & 0 & |0\end{matrix}\right]\]

ganeshie8 (ganeshie8):

Looks good! so above reduced system is equivalent to : ``` 5x + 3y - 9z = 2 - 8y - 76z = 33 ```

ganeshie8 (ganeshie8):

let \(z=0\) and find a particular solution

OpenStudy (clarence):

x-y=7?

ganeshie8 (ganeshie8):

how ? plugin \(z=0\) in the bottom equation first

OpenStudy (clarence):

Oh, I just plugged 0 in for z for both equations, my bad.

ganeshie8 (ganeshie8):

the advantage of triangular matrix is just that solving is trivial from bottom

OpenStudy (clarence):

So y = -33/8 for the bottom solution?

ganeshie8 (ganeshie8):

yes, find x also

ganeshie8 (ganeshie8):

plugin z=0, y=-33/8 in first equation

OpenStudy (clarence):

x = 23/8

ganeshie8 (ganeshie8):

right, so our particular solution is (23/8, -33/8, 0)

ganeshie8 (ganeshie8):

save that, we need to find nullspace

ganeshie8 (ganeshie8):

for nullspace, you want to solve : ``` 5x + 3y - 9z = 0 - 8y - 76z = 0 ```

ganeshie8 (ganeshie8):

plugin \(z=1\) and solve y, x values

OpenStudy (clarence):

x = 15/2 & y = -19/2?

ganeshie8 (ganeshie8):

Yes, so the nullspace is `t(15/2, -19/2, 1)`

ganeshie8 (ganeshie8):

the complete solution is given by `particular` + `nullspace`

OpenStudy (clarence):

So the one with z being 0 + what we just got?

ganeshie8 (ganeshie8):

Complete solution : ``` (23/8, -33/8, 0) + t(15/2, -19/2, 1) ```

OpenStudy (clarence):

Oh I see, that makes sense!